# Weak instrumental variables due to nonlinearities in panel data: A Super Learner Control Function estimator

### v2

@@ Abstract
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders with some of them time-invariant. In this setup, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to weak instruments As a solution, we propose a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. First, we estimate the control function using a super learner. In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the Super Learner Control Function Estimators significantly outperform Within 2SLS estimators.
Keywords— Endogeneity, Super Learner, Identification, Control Function, Panel data
@@ 1 Introduction
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders, and one some of them are time-invariant. In this setup, it is common to assume that the structural is linear or to exploit a linear projection in the reduced form equation. However, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to problems due to weak instruments since the strength of the linear relationship might not be high. This paper proposes to exploit the nonlinearties in the relationship between the endogenous covariate and the instrumental variables as a way to avoid problems of weak instruments due to nonlinearities.  [^1] Instruments are weak when they are weakly correlated to the endogenous covariate. In a cross-sectional setup, weak IVs causes that the median of the 2SLS estimator is biased towards the OLS estimator, 2SLS t-test presents size inflation (Keane and Neal (2024)). In panel data, Within-2SLS estimator is consistent when the time dimension grows (Cai et al. (2012)).
@@
$y_{it}=x_{1it}\beta_{1_{o}}+\alpha_{i,y}+\varepsilon_{it},\quad with\quad E[\varepsilon_{it}|z_{i},\alpha_{i,y}]=0,$ (1)
$x_{1it}=g_{o}(z_{it})+\alpha_{i,1x}+u_{it},\quad with\quad E[u_{it}|z_{i},\alpha_{i,1x}]=0,$ (2)
$\varepsilon_{it}=\rho u_{it}+\omega_{it},\quad with\quad E[\omega_{it}|u_{i},z_{i},\alpha_{i,y}]=0,$ (3)
@@
In Section 3, we present motivating examples. In Section 4, we present the baseline model. In Section 5 we describe the identification strategy. In Section 6, we describe the control function approach. In Section 7, we present the simulation experiment and its results, in Section 8 we present the results of the empirical application, and in Section 9 we present the conclusions of the paper.
@@ 3.1 Air pollution and educational outcomes
with $WD_{dt}$ is the average wind direction in district $d$ in year $t$, and $\gamma_{d}$ district specific effects.
@@ 3.2 Air pollution and health outcomes
$PM2.5_{cdmy}=h(\theta_{cdmy}),X_{cdmy}^{\prime})+\alpha_{c}+\alpha_{sm}+\alpha_{my}+\varepsilon_{cd},$ (9)
@@ 4 The Setup
We study a model composed of a linear structural equation, the outcome variable ($y_{it}$) explained by an endogenous variable ($x_{1it}$), $K$ exogenous regressors ($\tilde{x}_{it}$), and an additive individual-specific effect ($\alpha_{i,y}$). The model presents an external instrumental variable ($z_{it}$) for the endogenous regressor. The endogenous regressor is mean dependent on the exogenous regressors, the instrumental variable, and the individual specific effects $\alpha_{i,x}$. We assume that:
@@
We observe a sequence of data sets $\{y_{i},x_{1i},...,x_{Ki},z_{i}\}_{i=1}^{N}$, which are $T\times 1$ identical and independent copies of the $K+2$ random vectors $(\bm{y},\bm{x}_{1},\bm{x}_{2},...,\bm{x}_{K},\bm{z})$ taking values in $[-M_{y},M_{y}]^{T}\times[-M_{x1},M_{x1}]^{T}\times[-M_{xK},M_{xK}]^{T}\times[-M_{z},M_{z}]^{T}$ with $M_{y}<\infty$, $M_{xk}<\infty$, $M_{z}<\infty$. Thus, each element $t$ of $\bm{y}$ is bounded and contained in $[-M_{y},M_{y}]$. Similarly, each element of $x_{k}$ is contained in $[-M_{k},M_{k}]$, which means that $\bm{x}_{kt}$ is a continuously distributed random variable with compact support $[-M_{k},M_{k}]$. In addition, each element of $\bm{z}$ is a continuously distributed random variable with compact support $[-M_{z},M_{z}]$.
@@
$\begin{gathered}y_{it}=x_{1it}\beta_{1_{o}}+\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\alpha_{i,y}+\varepsilon_{it},\quad i\in\{1,...,N\},t\in\{1,...,T\},\\ E[\varepsilon_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y}]=0.\end{gathered}$ (10)
@@
The unobserved random term $\varepsilon_{it}$ is independent and identically distributed with a continuous distribution. It has a compact support with zero mean conditional on $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT}]^{\prime}$, $\alpha_{i,y}$, and $Var(\varepsilon_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y})<\infty$. The unobserved parameter of interest is $\beta_{1_{o}}$.
@@
$\begin{gathered}x_{1it}=g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,1x}+u_{it},\quad i\in\{1,...,N\},t\in\{1,...,T\},\\ E[u_{it}|\tilde{x}_{i},z_{i},\alpha_{i,1x}]=0.\end{gathered}$ (11)
with the unknown function $g_{o}:[-M_{x},M_{x}]^{K}\times[-M_{z},M_{z}]\rightarrow\mathcal{G}\subset\mathbb{R}$.
The reduced form error term $u_{it}$ is independent and identically distributed with a continuous distribution. It has a compact support with zero mean conditional on $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT}]^{\prime}$, $\alpha_{i,x}$, and $Var(u_{it}|\tilde{x}_{i},z_{i},\alpha_{i,1x})<\infty$.
@@
$\begin{gathered}\varepsilon_{it}=\rho u_{it}+\omega_{it},\quad i\in\{1,...,N\},t\in\{1,...,T\},\\ E[\omega_{it}|u_{i},\tilde{x}_{i},z_{i},\alpha_{i,y}]=0.\end{gathered}$ (12)
Assumption 4 states that $\omega_{it}$ is strictly exogenous of $u_{it}$. The error term $\omega_{it}$ is is independent and identically distributed with a continuous distribution. It has a bounded compact support with zero mean conditional on $u_{i}=[u_{i1},u_{i2},,...,u_{iT}]^{\prime}$, $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT}]^{\prime}$, $\alpha_{i,y}$, and $Var(\omega_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y})<\infty$.
As consequence of Assumptions 3 and 4, the structural error term $\varepsilon_{it}$ is correlated with $x_{1it}$. The reason is that $E[x_{1it}\varepsilon_{it}]=E[(g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,x}+u_{it})\varepsilon_{it}]$ by Assumption 3. The last expression is equal to $E[(g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,x}+u_{it})(\rho_{o}u_{it}+\omega_{it})]=\rho_{o}E[u_{it}^{2}]$ by Assumption 4.
The error term $\omega_{it}$ has zero mean conditional on $x_{1it}$, $\tilde{x}_{it}$, $\alpha_{i,y}$, and $u_{it}$ ($E[\omega_{it}|x_{1it},\tilde{x}_{it},\alpha_{i,y}]=0$) because $E[\omega_{it}|x_{1it},\tilde{x}_{it},u_{it},\alpha_{i,y}]=E[\omega_{it}|\tilde{x}_{it},z_{it},\alpha_{i,y},u_{it}]=0$ where the first equality holds because $x_{1it}$ is a one-to-one function of $z_{it}$ conditional on $\tilde{x}_{it}$, $u_{it}$, and $\alpha_{i,y}$ (Wooldridge, 2010). The second equality holds by Assumption 4.
@@ 5 Identification: Control Function Approach
The identification of the parameter of interest $\beta_{1_{o}}$ relies on the presence of an instrumental variable $z_{it}$ conditioning on the individual-specific effects (Assumptions 2 and 3). Since we run into an incidental parameter problem, we get rid of the individual-specific effects by using an appropriate transformation of the data $\mathbf{T}:\mathbb{R}\rightarrow\mathbb{R}$. The typical transformations of the data are first-differencing and within transformation. The first-difference transformation is equal to $y_{it}-y_{it-1}$, and the Within transformation is equal to $y_{it}-\sum_{t=1}^{T}y_{it}/T$.
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\mathbf{T}\varepsilon_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\},$ (13)
$\mathbf{T}x_{1it}=\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})+\mathbf{T}u_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\}.$ (14)
@@
$\mathbf{T}y_{it}=\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\mathbf{T}u_{it}\beta_{1_{o}}+\mathbf{T}\varepsilon_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\}.$ (15)
In equation 15, one can believe that the estimation of the unknown parameters is straightforward since the error term $\mathbf{T}u_{it}\beta_{1_{o}}+\mathbf{T}\varepsilon_{it}$ in 15 has zero mean conditional $\tilde{x}_{it-1}$, $\tilde{x}_{it-1}$, $z_{it}$, and $z_{it-1}$ as a consequence of assumptions 2, and 3. Moreover, one can think that the best estimation procedure is to estimate $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})$ in the first stage and in the second stage, plug it in equation 15 and perform simple OLS. This is not possible because the regularization bias of the first stage estimator of $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})$ contaminates the estimation of $\beta_{1_{o}}$ and $\beta_{2_{o}}$ (Chernozhukov et al., 2018; Robinson, 1988; Guo et al., 2022) and produces estimates that are not $\sqrt{N}$-consistent.
@@
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\rho_{o}\mathbf{T}u_{it}+\mathbf{T}\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T\}.$ (16)
In model 16, we can guarantee zero correlation between $\mathbf{T}x_{1it}$, and $\mathbf{T}u_{it}$ thanks to the presence of an instrumental variable in the reduced form equation. In addition, we can set up the following moment conditions:
$E[\phi(\mathbf{T}H_{it};\theta_{o},\mathbf{T}g_{o}(\tilde{x}_{it},z_{it}))]=\textbf{0},$ (17)
where $\phi(\mathbf{T}H_{it};\theta_{o},\mathbf{T}g_{o}(\tilde{x}_{it},z_{it}))$ is a vector of score functions equal to $\mathbf{T}\omega_{it}\mathbf{T}H_{it}$, $\theta_{o}=[\beta_{1_{o}}\quad\beta_{2_{o}}^{\prime}\quad\rho_{o}]^{\prime}$, $\mathbf{T}\omega_{it}=\mathbf{T}y_{it}-\mathbf{T}x_{1it}\beta_{1_{o}}-\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}-\rho_{o}\mathbf{T}u_{it}$, 0=[$\textbf{0}=[0\quad 0\quad 0]^{\prime}$, and $\mathbf{T}H_{it}=[\mathbf{T}x_{1it}\quad\mathbf{T}\tilde{x}_{it}\quad\mathbf{T}u_{it}]^{\prime}.$
These moment conditions can be seen as the first-order conditions of the minimization of the following loss function:
$Q_{o}(\mathbf{T}H_{it};\theta,\mathbf{T}g_{o}(\tilde{x}_{it},z_{it}))=E[(\mathbf{T}y_{it}-\mathbf{T}x_{1it}\beta_{1}-\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k}-\mathbf{T}u_{it}\rho)^{2}].$
This interpretation is important for consistency of the estimator proposed in Section 6 because it permits to specify the necessary primitive condition that $Q_{o}(\mathbf{T}H_{it};\theta,\mathbf{T}g_{o}(\tilde{x}_{it},z_{it}))$ has a unique global optimizer $\theta_{o}$ (Newey and McFadden, 1994).
@@
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\Delta\varepsilon_{it},i\in\{1,...,N\},t\in\{2,...,T\}.$ (18)
$\Delta x_{1it}=\Delta g_{o}(\tilde{x}_{it},z_{it})+\Delta u_{it},i\in\{1,...,N\},t\in\{2,...,T\}.$ (19)
@@
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\Delta u_{it}+\Delta\omega_{it},i\in\{1,...,N\},t\in\{2,...,T\}.$ (20)
@@
$E[\Delta\omega_{it}\Delta H_{it}]=\textbf{0},$ (21)
where $\Delta\omega_{it}=\Delta y_{it}-\Delta x_{1it}\beta_{1_{o}}-\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}-\rho_{o}\Delta u_{it}$, 0=[$\textbf{0}=[0\quad 0\quad 0]^{\prime}$, and $\Delta H_{it}=[\Delta x_{1it}\quad\Delta\tilde{x}_{it}\quad\Delta u_{it}]^{\prime}.$
@@
Under assumption 3, $\Delta g({\tilde{x}_{it},z_{it}})$ is identified by $E[\Delta x_{1it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]$.
@@
$E[\Delta x_{1it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=E[\Delta g(\tilde{x}_{it},z_{it})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]+E[\Delta u_{it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]$ (22)

### v3

@@ Abstract
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders with some of them time-invariant. In this setup, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to weak instruments (instruments are weakly correlated with the endogeneous covariate). As a solution, we propose a triangular simultaneous equation model for panel data with additive separable individual-specific fixed effects composed of a linear structural equation with a nonlinear reduced form equation. The parameter of interest is the structural parameter of the endogenous variable. The identification of this parameter is obtained under the assumption of available exclusion restrictions and using a control function approach. Estimating the parameter of interest is done using an estimator that we call Super Learner Control Function estimator (SLCFE). The estimation procedure is composed of two main steps and sample splitting. First, we estimate the control function using a super learner . In the following step, we use the estimated control function to control for endogeneity in the structural equation. Sample splitting is done across the individual dimension. The estimator is consistent. We perform a Monte Carlo simulation to test the performance of the estimators proposed. We conclude that the Super Learner Control Function Estimators significantly outperform Within 2SLS estimators.
Keywords— Endogeneity, Super Learner, Identification, Control Function, Panel data, Individual Fixed Effects
@@ 1 Introduction
A triangular structural panel data model with additive separable individual-specific effects is used to model the causal effect of a covariate on an outcome variable when there are unobservable confounders, and some of them are time-invariant. In this setup, it is common to assume that the structural is linear or to exploit a linear projection in the reduced form equation. However, a linear reduced-form equation might be problematic when the conditional mean of the endogenous covariate and the instrumental variables is nonlinear. The reason is that ignoring the nonlinearity could lead to problems due to weak instruments since the strength of the linear relationship might not be high. This paper proposes to exploit the nonlinearties in the relationship between the endogenous covariate and the instrumental variables as a way to avoid problems of weak instruments due to nonlinearities.  [^1] Instruments are weak when they are weakly correlated to the endogenous covariate. In a cross-sectional setup, weak IVs causes that the median of the 2SLS estimator is biased towards the OLS estimator, 2SLS t-test presents size inflation (Keane and Neal (2024)). In panel data, Within-2SLS estimator is consistent when the time dimension grows (Cai et al. (2012)).
@@
$y_{it}=x_{1it}\beta_{1_{o}}+\alpha_{i,y}+\varepsilon_{it},\quad with\quad\mathbb{E}[\varepsilon_{it}|z_{i},\alpha_{i,y}]=0,$ (1)
$x_{1it}=g_{o}(z_{it})+\alpha_{i,1x}+u_{it},\quad with\quad\mathbb{E}[u_{it}|z_{i},\alpha_{i,1x}]=0,$ (2)
$\varepsilon_{it}=\rho u_{it}+\omega_{it},\quad with\quad\mathbb{E}[\omega_{it}|u_{i},z_{i},\alpha_{i,y}]=0,$ (3)
@@
In Section 3, we present motivating examples. In Section 4, we present the baseline model. In Section 5 we describe the identification strategy. In Section 6, we describe the control function approach. In Section 8, we present the simulation experiment and its results, in Section 7 we present the large sample properties of the SLCF estimator, in Section 9 we present the results of the empirical application, and in Section 10 we present the conclusions of the paper.
@@ 3.1 Air pollution and educational outcomes
where $WD_{dt}$ is the average wind direction in district $d$ in year $t$, and $\gamma_{d}$ represent district specific effects.
@@ 3.2 Air pollution and health outcomes
$PM2.5_{cdmy}=h(\theta_{cdmy},X_{cdmy}^{\prime})+\alpha_{c}+\alpha_{sm}+\alpha_{my}+\varepsilon_{cd},$ (9)
@@ 4 The Setup
We study a model composed of a linear structural equation, the outcome variable ($y_{it}$) explained by an endogenous variable ($x_{1it}$), $K-1$ exogenous regressors ($\tilde{x}_{it}$), and an additive individual-specific effect ($\alpha_{i,y}$). The model presents an external instrumental variable ($z_{it}$) for the endogenous regressor. The endogenous regressor is mean dependent on the exogenous regressors, the instrumental variable, and the individual specific effects $\alpha_{i,x}$. We assume that:
@@
We observe a sequence of data sets $\{O_{i}=(y_{i},x_{1i},...,x_{Ki},z_{i})\}_{i=1}^{N}$, which are $T_{i}\times 1$ independent copies of the $K+2$ random vectors $(\mathbf{y},\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{K},\mathbf{z})$ taking values in $[-M_{y},M_{y}]^{T}\times[-M_{x1},M_{x1}]^{T}\times[-M_{xK},M_{xK}]^{T}\times[-M_{z},M_{z}]^{T}$ with $M_{y}<\infty$, $M_{xk}<\infty$, $M_{z}<\infty$. Thus, each element $t$ of $\mathbf{y}$ is bounded and contained in $[-M_{y},M_{y}]$. Similarly, each element of $x_{k}$ is contained in $[-M_{k},M_{k}]$, which means that $\mathbf{x}_{kt}$ is a continuously distributed random variable with compact support $[-M_{k},M_{k}]$. In addition, each element of $\mathbf{z}$ is a continuously distributed random variable with compact support $[-M_{z},M_{z}]$.
Assumption 1 states that each individual $i$ presents $T_{i}$ observations. We assume that $T_{i}$ is uniformly bounded such that $T_{i}<T_{max}$. The total number of observations is denoted by $N_{T}=\sum_{i}T_{i}$.
@@
$\begin{gathered}y_{it}=x_{1it}\beta_{1_{o}}+\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\alpha_{i,y}+\varepsilon_{it},\quad i\in\{1,...,N\},t\in\{1,...,T_{i}\},\\ \mathbb{E}[\varepsilon_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y}]=0.\end{gathered}$ (10)
@@
The unobserved random term $\varepsilon_{it}$ is independent and identically distributed with a continuous distribution. It has a compact support with zero mean conditional on $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT_{i}}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT_{i}}]^{\prime}$, $\alpha_{i,y}$, and $Var(\varepsilon_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y})<\infty$. The unobserved parameter of interest is $\beta_{1_{o}}$.
@@
$\begin{gathered}x_{1it}=g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,1x}+u_{it},\quad i\in\{1,...,N\},t\in\{1,...,T_{i}\},\\ \mathbb{E}[u_{it}|\tilde{x}_{i},z_{i},\alpha_{i,1x}]=0.\end{gathered}$ (11)
with the unknown function $g_{o}:[-M_{x},M_{x}]^{K}\times[-M_{z},M_{z}]\rightarrow\mathcal{S}\subset\mathbb{R}$.
The reduced form error term $u_{it}$ is independent and identically distributed with a continuous distribution. It has a compact support with zero mean conditional on $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT_{i}}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT_{i}}]^{\prime}$, $\alpha_{i,x}$, and $Var(u_{it}|\tilde{x}_{i},z_{i},\alpha_{i,1x})<\infty$.
@@
$\begin{gathered}\varepsilon_{it}=\rho u_{it}+\omega_{it},\quad i\in\{1,...,N\},t\in\{1,...,T_{i}\},\\ \mathbb{E}[\omega_{it}|u_{i},\tilde{x}_{i},z_{i},\alpha_{i,y}]=0.\end{gathered}$ (12)
Assumption 4 states that $\omega_{it}$ is strictly exogenous of $u_{it}$. The error term $\omega_{it}$ is is independent and identically distributed with a continuous distribution. It has a bounded compact support with zero mean conditional on $u_{i}=[u_{i1},u_{i2},,...,u_{iT_{i}}]^{\prime}$, $\tilde{x}_{i}=[\tilde{x}_{i1},\tilde{x}_{i2},,...,\tilde{x}_{iT_{i}}]^{\prime}$, $z_{i}=[z_{i1},z_{i2},...,z_{iT_{i}}]^{\prime}$, $\alpha_{i,y}$, and $Var(\omega_{it}|\tilde{x}_{i},z_{i},\alpha_{i,y})=\sigma^{2}_{\omega}<\infty$. This assumption is relaxed in a follow up study.
As consequence of Assumptions 3 and 4, the structural error term $\varepsilon_{it}$ is correlated with $x_{1it}$. The reason is that $\mathbb{E}[x_{1it}\varepsilon_{it}]=\mathbb{E}[(g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,x}+u_{it})\varepsilon_{it}]$ by Assumption 3. The last expression is equal to $\mathbb{E}[(g_{o}(\tilde{x}_{it},z_{it})+\alpha_{i,x}+u_{it})(\rho_{o}u_{it}+\omega_{it})]=\rho_{o}\mathbb{E}[u_{it}^{2}]$ by Assumption 4.
The error term $\omega_{it}$ has zero mean conditional on $x_{1it}$, $\tilde{x}_{it}$, $\alpha_{i,y}$, and $u_{it}$ ($\mathbb{E}[\omega_{it}|x_{1it},\tilde{x}_{it},u_{it},\alpha_{i,y}]=0$) because $\mathbb{E}[\omega_{it}|x_{1it},\tilde{x}_{it},u_{it},\alpha_{i,y}]=\mathbb{E}[\omega_{it}|\tilde{x}_{it},z_{it},\alpha_{i,y},u_{it}]=0$ where the first equality holds because $x_{1it}$ is a one-to-one function of $z_{it}$ conditional on $\tilde{x}_{it}$, $u_{it}$, and $\alpha_{i,y}$ (Wooldridge, 2010). The second equality holds by Assumption 4.
@@ 5 Identification: Control Function Approach
The identification of the parameter of interest $\beta_{1_{o}}$ relies on the presence of an instrumental variable $z_{it}$ conditioning on the individual-specific effects (Assumptions 2 and 3). Since we run into an incidental parameter problem, we get rid of the individual-specific effects by using an appropriate transformation of the data $\mathbf{T}:\mathbb{R}\rightarrow\mathbb{R}$. The typical transformations of the data are first-differencing and within transformation. The first-difference transformation is equal to $y_{it}-y_{it-1}$, and the Within transformation is equal to $y_{it}-\sum_{t=1}^{T_{i}}y_{it}/T_{i}$.
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\mathbf{T}\varepsilon_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\},$ (13)
$\mathbf{T}x_{1it}=\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})+\mathbf{T}u_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (14)
@@
$\mathbf{T}y_{it}=\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\mathbf{T}u_{it}\beta_{1_{o}}+\mathbf{T}\varepsilon_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (15)
In equation 15, one can believe that the estimation of the unknown parameters is straightforward since the error term $\mathbf{T}u_{it}\beta_{1_{o}}+\mathbf{T}\varepsilon_{it}$ in 15 has zero mean conditional an appropriate information set $I_{t}=\{\tilde{x}_{it-jk},z_{it-j},0\leq j\leq A\}$ with $A\leq t-1$ as a consequence of assumptions 2, and 3. Moreover, one can think that the best estimation procedure is to estimate $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})$ in the first stage and in the second stage, plug it in equation 15 and perform simple OLS. This is not possible because the regularization bias of the first stage estimator of $\mathbf{T}g_{o}(\tilde{x}_{it},z_{it})$ contaminates the estimation of $\beta_{1_{o}}$ and $\beta_{2_{o}}$ (Chernozhukov et al., 2018; Robinson, 1988; Guo et al., 2022) and produces estimates that are not $\sqrt{N}$-consistent.
@@
$\mathbf{T}y_{it}=\mathbf{T}x_{1it}\beta_{1_{o}}+\sum_{k=2}^{K}\mathbf{T}x_{kit}\beta_{2k_{o}}+\rho_{o}\mathbf{T}u_{it}+\mathbf{T}\omega_{it},i\in\{1,...,N\},t\in\{t_{a},...,T_{i}\}.$ (16)
In model 16, we can guarantee zero correlation between $\mathbf{T}x_{1it}$, and $\mathbf{T}u_{it}$ thanks to the presence of an instrumental variable in the reduced form equation.
If we stack up the time observations and denote as $\mathbb{T}_{i}$ the transformation operator that maps $\mathbb{R}^{T_{i}}\rightarrow\mathbb{R}^{T_{i,a}}$ with $T_{i,a}$ equal to $T$ if we use the within transformation and $T-1$ if we use the first-difference transformation, we set up the following moment conditions  [^3] We acknowledge that $\mathbb{T}$ traditionally denotes the circle group in mathematics; however, we adopt similar notation here for convenience.:
$\mathbb{E}[\phi(O_{i};\theta_{o},\mathbf{T}g_{o}]=\textbf{0},$ (17)
where $\phi(O_{i};\theta_{o},\mathbf{T}g_{o})$ is a vector of score functions equal to $(\mathbb{T}_{i}H_{i})^{\prime}V_{i}^{-1}\mathbb{T}_{i}\omega_{i}$, $\theta_{o}=[\beta_{1_{o}}\quad\beta_{2_{o}}^{\prime}\quad\rho_{o}]^{\prime}$, $\mathbb{T}_{i}\omega_{i}=\mathbb{T}_{i}y_{i}-\mathbb{T}_{i}x_{1i}\beta_{1_{o}}-\sum_{k=2}^{K}\mathbb{T}_{i}x_{ki}\beta_{2k_{o}}-\rho_{o}\mathbb{T}_{i}u_{i}$, 0=[$\textbf{0}=[0\quad 0\quad 0]^{\prime}$, and $\mathbb{T}_{i}H_{i}=[\mathbb{T}_{i}x_{1i}\quad\mathbb{T}_{i}\tilde{x}_{i}\quad\mathbb{T}_{i}u_{i}]$, and $V_{i}=\mathbb{E}[(\mathbb{T}_{i}\omega_{i})(\mathbb{T}_{i}\omega_{i})^{\prime}|H_{i},z_{i}]=\sigma^{2}_{\omega}\mathbb{T}_{i}\mathbb{T}_{i}^{\prime}$.
Remark 1.
If we use the first-difference transformation, $\mathbb{T}_{i}=D_{i}=\begin{bmatrix}1&-1&0&\cdots&0\\ 0&1&-1&\cdots&0\\ \vdots&&\ddots&\ddots&\vdots\\ 0&\cdots&0&1&-1\\ \end{bmatrix}$, with dimensions $(T_{i}-1)\times T_{i}$. If we use the within transformation, $\mathbb{T}_{i}=W_{i}=1/T_{i}\begin{bmatrix}1&1&1&\cdots&1\\ 1&1&1&\cdots&1\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 1&\cdots&1&1&1\\ \end{bmatrix}$ , with dimensions $T_{i}\times T_{i}$.
Remark 2.
The moment condition is valid even under misspecification of $V_{i}$ because we assume strict exogeneity. If we would like to relax the assumption of strict exogeneity to weak exogeneity, then it is better to use $\mathbb{E}[(\mathbb{T}_{i}H_{i})^{\prime}\mathbb{T}_{i}\omega_{i}]=0$ as moment conditions.
Remark 3.
If $\mathbb{T}_{i}=W_{i}$ (the within transformation), the moment condition can be simply written as $\mathbb{E}[\mathbb{T}_{i}\omega_{i}^{\prime}\mathbb{T}_{i}H_{i}]=0$.
@@
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\Delta\varepsilon_{it},i\in\{1,...,N\},t\in\{2,...,T_{i}\}.$ (18)
$\Delta x_{1it}=\Delta g_{o}(\tilde{x}_{it},z_{it})+\Delta u_{it},i\in\{1,...,N\},t\in\{2,...,T_{i}\}.$ (19)
@@
$\Delta y_{it}=\Delta x_{1it}\beta_{1_{o}}+\Delta\tilde{x}_{it}^{\prime}\beta_{2_{o}}+\rho_{o}\Delta u_{it}+\Delta\omega_{it},i\in\{1,...,N\},t\in\{2,...,T_{i}\}.$ (20)
@@
$\mathbb{E}[(D_{i}H_{i})^{\prime}V_{i}^{-1}D_{i}\omega_{i}]=\textbf{0},$ (21)
where $D_{i}\omega_{i}=D_{i}y_{i}-D_{i}x_{1i}\beta_{1_{o}}-D_{i}\tilde{x}_{i}^{\prime}\beta_{2_{o}}-\rho_{o}D_{i}u_{i}$, 0=[$\textbf{0}=[0\quad 0\quad 0]^{\prime}$, $D_{i}H_{i}=[D_{i}x_{1i}\quad D_{i}\tilde{x}_{i}\quad D_{i}u_{i}]^{\prime}$, and $V_{i}=\sigma^{2}_{\omega}D_{i}D_{i}^{\prime}$.
@@
Under assumption 3, $\Delta g({\tilde{x}_{it},z_{it}})$ is identified by $\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]$.
@@
$\mathbb{E}[\Delta x_{1it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]=\mathbb{E}[\Delta g(\tilde{x}_{it},z_{it})|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]+\mathbb{E}[\Delta u_{it}|\tilde{x}_{it},z_{it},\tilde{x}_{it-1},z_{it-1}]$ (22)